OpenStudy (anonymous):

Evaluate: [(21 + 6) − 3*2] ÷ 9 • 2 4 4.6 1 52

OpenStudy (anonymous):

is the 2 on the denominator? If so then it becomes 21/18 = 1.666 which doesnt seem to be one of the options...there might be a typo

OpenStudy (johnweldon1993):

I think that is how it's written purely for an order of operations question [(21 + 6) − 3*2] ÷ 9 • 2 deal with the innermost parenthesis first [(27) − 3*2] ÷ 9 • 2 now whats left in the brackets [(27) − 6] ÷ 9 • 2 Finish up the brackets 21 ÷ 9 • 2 Now, since you only have multiplication and division left....go left to right 21 divided by 9 = 2.3 That times 2 = 4.6 Which is one of your choices! Hope that helped!

OpenStudy (anonymous):

the 2 is a exponint

OpenStudy (johnweldon1993):

So it should be [(21 + 6) − 3²] ÷ 9 • 2 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then the answer is 4

OpenStudy (johnweldon1993):

If so then it comes out to (27 - 9) = 18 now 18 ÷ 9 • 2 Left to right 18÷ 9 = 2 2• 2 = 4

OpenStudy (anonymous):

Peter's age can be described as 2 more than 3 times his son's age. If his son's age is represented by the variable "s", write the algebraic expression representing Peter's age using "s".

OpenStudy (johnweldon1993):

Well Peter is 2 more ( +2 ) than 3 times his sons age (3s) so we can say that P = 3s + 2